The object of mass m shown above is dropped from rest near Earth’s surface and experiences a resistive force of magnitude kv , where v is the speed of the object and k is a constant. Which of the following expressions can be used to find v as a function of time t ? (Assume that the direction of the gravitational force is positive.)
1
Expert's answer
2019-11-06T09:20:48-0500
"V = \frac{m(g+V_{o} /t)}{k+m/t}"


Where V - speed of object, m - mass of object, Vo -initial speed of object, t - time,


"mg-kV=ma"

Where a - acceleration of object


"a= \frac{V-V_{o}}{t}"

"mg = kV + m(\frac{V-V_{o}}{t})"

"V = \frac{m(g+V_{o} /t)}{k+m/t}"